# Discovering and exploiting active sensing motifs for estimation

### v1

@@ Methods
This work was supported by a National Science Postdoctoral Fellowship in Biology to Benjamin Cellini (NSF 22-623). The work was also supported by the Air Force Office of Scientific Research (FA9550-21-0122) to FvB, the NSF AI institute in Dynamics (2112085) to FvB, NSF-EFRI-BRAID (2318081) to FvB, Sloan Research Fellowship (FG-2020-13422) to FvB, and DEPSCoR (FA9950-23-1-0483) to FvB. This research was also supported in part by the NSF grant PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), which supported valuable discussions that led to substantial improvements of the paper. For the purpose of open access, the author has applied a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission.
@@ Acknowledgments
We thank Christina May for productive conversations about applications of observability analysis on real data and feedback on pybounds code, Stanley David Stupski for providing data for testing, and J. Humberto Ramos for help developing the quadcopter dynamical model.
@@
Conceptualization: B.C, B.B, F.v.B.; Methodology: B.C, B.B, A.L, F.v.B.; Software: B.C; Formal analysis: B.C; Investigation: B.C, F.v.B.; Resources: F.v.B.; Data curation: B.C, A.L; Writing - original draft: B.C, F.v.B.; Writing - review & editing: B.C, B.B, A.L, F.v.B.; Visualization: B.C, F.v.B.; Supervision: F.v.B.; Project administration: F.v.B.; Funding acquisition: B.C, F.v.B.;
@@
Supplementary Figure S4: ANN Wind direction estimator performance. Wind direction artificial neural network (ANN) estimator (see LABEL:fig:figure_3_c) testing performance as a function of training and testing data with different observability levels. The dataset in LABEL:fig:figure_3_b was split into 10 equally sized bins and an ANN was trained on 80% of the data from each bin. The withheld 20% of each bin was used to evaluate the performance. LABEL:fig:figure_3_c shows the testing estimation error variance $\text{Var}(\zeta-\check{\zeta})$ for each bin corresponding to each panel here compared to ANN’s trained on randomly split data. The ANN input (measurement) noise variance were all set to $10^{-4}$ rad, corresponding to a noise covariance of $\mathcal{R}=10^{-4}\times I_{p\times p}$.
@@
Supplementary Figure S6: ANN altitude estimator. a. Artificial neural network (ANN) that estimates altitude from measurements of forward optic flow ($r_{x}$) and forward acceleration ($\dot{v}_{x}$) over the time widow $\omega$. b. Altitude ANN estimator testing performance as a function of the input (measurement) noise levels, where $\mathcal{R}=10^{-4}$ indicates a noise covariance matrix of $\mathcal{R}=10^{-4}\times I_{p\times p}$ for the ANN inputs. The gray dashed line indicates perfect performance. c. Illustration of mapping the mean forward acceleration ($\dot{v}_{x}$) in the window $\omega$ to the augmented noise variance $\check{\mathcal{R}}_{z,z}$ for the altitude estimate. d. Altitude ANN estimator testing performance as a function of the mean forward acceleration ($\dot{v}_{x}$) in the window $\omega$, as a function of the input (measurement) noise levels. Note that the shape of the heatmap illustrates the variance of the altitude estimates.

### v2

@@ 7. Reveal active sensing motifs by correlating state dynamics with observability
This work was supported by a National Science Postdoctoral Fellowship in Biology to Benjamin Cellini (NSF 22-623). The work was also supported by the Air Force Office of Scientific Research (FA9550-21-0122) to FvB, the NSF AI institute in Dynamics (2112085) to FvB, NSF-EFRI-BRAID (2318081) to FvB, Sloan Research Fellowship (FG-2020-13422) to FvB, DEPSCoR (FA9950-23-1-0483) to FvB, and NIH (1R01NS136988) to FvB. This research was also supported in part by the NSF grant PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), which supported valuable discussions that led to substantial improvements of the paper. For the purpose of open access, the author has applied a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission.
@@ Acknowledgments
We thank Christina May for productive conversations about applications of observability analysis on real data and feedback on pybounds code, Stanley David Stupski for providing data for testing, J. Humberto Ramos for help developing the quadcopter dynamical model, and Natalie Brace and Aditya Nair for helpful comments on the manuscript.
@@
Conceptualization: B.C., B.B., F.v.B.; Methodology: B.C., B.B., A.L., F.v.B.; Software: B.C.; Formal analysis: B.C.; Investigation: B.C., F.v.B.; Resources: F.v.B.; Data curation: B.C., A.L.; Writing - original draft: B.C., F.v.B.; Writing - review & editing: B.C., B.B., A.L., F.v.B.; Visualization: B.C., F.v.B.; Supervision: F.v.B.; Project administration: F.v.B.; Funding acquisition: B.C., F.v.B.
@@
Supplementary Figure S4: ANN Wind direction estimator performance. Wind direction artificial neural network (ANN) estimator (see LABEL:fig:figure_3_c) testing performance as a function of training and testing data with different observability levels. The dataset in LABEL:fig:figure_3_b was split into 10 equally sized bins and an ANN was trained on 80% of the data from each bin. The withheld 20% of each bin was used to evaluate the performance. LABEL:fig:figure_3_c shows the testing estimation error variance $\text{Var}(\zeta-\check{\zeta})$ for each bin corresponding to each panel here compared to ANNs trained on randomly split data. The ANN input (measurement) noise variance were all set to $10^{-4}$ rad, corresponding to a noise covariance of $\mathcal{R}=10^{-4}\times I_{p\times p}$.
@@
Supplementary Figure S6: ANN altitude estimator. a. Artificial neural network (ANN) that estimates altitude from measurements of forward optic flow ($r_{x}$) and forward acceleration ($\dot{v}_{x}$) over the time window $\omega$. b. Altitude ANN estimator testing performance as a function of the input (measurement) noise levels, where $\mathcal{R}=10^{-4}$ indicates a noise covariance matrix of $\mathcal{R}=10^{-4}I_{p\times p}$ for the ANN inputs. The gray dashed line indicates perfect performance. c. Illustration of mapping the mean forward acceleration ($\dot{v}_{x}$) in the window $\omega$ to the augmented noise variance $\check{\mathcal{R}}_{z,z}$ for the altitude estimate. d. Altitude ANN estimator testing performance as a function of the mean forward acceleration ($\dot{v}_{x}$) in the window $\omega$, as a function of the input (measurement) noise levels. Note that the shape of the heatmap illustrates the variance of the altitude estimates.